# Advancing Multimodal Reasoning: From Optimized Cold Start to Staged Reinforcement Learning

### v1

@@ 6.1 Multimodal Large Language Model
Sophisticated reasoning in Multimodal Large Language Models (MLLMs), inspired by Large Language Model (LLM) advancements, is a key research area. Initial progress involved integrating visual encoders with LLMs (e.g., CLIP radford2021learning , MiniGPT4 zhu2024minigpt ), with notable advancements driven by visual instruction tuning in series like LLaVA li2024llava li2024llavanextinterleavetacklingmultiimagevideo . While leading closed-source models (e.g., GPT-o3 chatgpto3 , Kimi-VL team2025kimi ) excel at long Chain-of-Thought (CoT) reasoning, open-source contributions have focused on CoT adaptations liu2023visualinstructiontuning ; guo2024mammothvlelicitingmultimodalreasoning ; su2024timo and Supervised Fine-Tuning (SFT) with reasoning traces mitra2024compositionalchainofthoughtpromptinglarge gao2024factteachingmllmsfaithful . The open-sourcing of DeepSeek-R1 guo2025deepseek has further spurred Reinforcement Learning (RL) applications for visual reasoning visionr1 dong2024insight and specialized domains like mathematical reasoning. Nevertheless, many MLLMs reasoning models wang2025vl openvlthinker r1-onevision chen2025r1v are limited by generating relatively short responses, which often curtail genuine reflection, thorough visual exploration, and consequently, deep multimodal reasoning. Our work, in contrast, introduces a novel framework to enable MLLMs to generate significantly longer, reflective responses with explicit visual references, thereby facilitating long CoT reasoning to unlock more comprehensive multimodal reasoning capabilities.
@@ 6.2 Reinforcement Learning in Reasoning
Reinforcement learning (RL) significantly advances LLM reasoning, evolving from methods like Reinforcement Learning from Human Feedback (RLHF) ouyang2022training ; li2025test . Current LLM research further explores direct RL fine-tuning, specialized cold-start datasets for long-form reasoning, and advanced algorithms like Group Relative Policy Optimization (GRPO) shao2024deepseekmath and its refinements (e.g., DAPO yu2025dapo , DR.GRPO liu2025understanding , GPG chu2025gpgsimplestrongreinforcement ) to elicit deeper reasoning. However, RL application to multimodal reasoning in MLLMs is nascent. Initial MLLMs efforts focus on subdomains like math reasoning meng2025mm ; peng2025lmm or generative reward models gao2025benchmarkingmultimodalcotreward , often utilizing data from commercial models. Nonetheless, successes such as DeepSeek-R1’s guo2025deepseek rule-based RL are spurring similar MLLMs investigations, indicating growing interest in RL for unlocking sophisticated multimodal reasoning.
@@ 7 Conclusion
This paper introduces ReVisual-R1, a 7B open-source MLLM designed to address prevalent challenges in cultivating sophisticated multimodal reasoning. By systematically integrating a strategic, high-difficulty text-only cold-start phase for foundational reasoning, a Multimodal RL stage employing GRPO stabilized by our novel Prioritized Advantage Distillation (PAD) mechanism and guided by rule-based rewards including an Efficient-Length Reward, and a final TextRL refinement phase, our structured three-stage curriculum demonstrates that thoughtful data strategy and targeted algorithmic optimizations are pivotal. ReVisual-R1 achieves state-of-the-art performance among open-source 7B models on a suite of challenging visuo-mathematical and reasoning benchmarks. This work underscores that careful curriculum design and algorithmic enhancements, rather than sheer model scale, can unlock robust, self-reflective multimodal reasoning.
Appendix A Appendix
A.1 Training settings
@@
Figure 5: Prompt Template for ReVisual-R1 in both inference and training stages.
A.2 Algorithm in Prioritized Advantage Distillation (PAD)
@@
Initially, PAD filters the original batch $\mathcal{B}$ to create an “effective set” $\mathcal{E}$ of sample indices and a corresponding map $\hat{A}_{\mathcal{E}}$ for their advantages (Lines 2-10 in Algorithm 1). For each sequence $i$ in $\mathcal{B}$, its absolute advantage $|\tilde{A}_{i}|$ is computed. If this value falls within a specified informative range $[T_{low},T_{high}]$, where $T_{low}>0$ is crucial for excluding stagnant (near-zero advantage) samples, the index $i$ is added to $\mathcal{E}$, and its absolute advantage $\hat{A}_{i,abs}$ is stored in $\hat{A}_{\mathcal{E}}$.
@@
1:Original batch =$\mathcal{B}=\{seq_{1},\dots,seq_{N}\}$ with $N=|\mathcal{B}|$; advantage estimates $A_{\mathrm{est}}=\{\tilde{A}_{1},\dots,\tilde{A}_{N}\}$; thresholds $T_{\mathrm{low}},\,T_{\mathrm{high}}$; temperature $\tau$; sub–sampling ratio $\rho$
2:Distilled mini-batch $\mathcal{B}_{\mathrm{distilled}}$
3$\triangleright$ Steps 1 & 2: Per-Sequence Advantage Calculation and Effective Sample Filtering
$\mathcal{E}\leftarrow\emptyset$triangleright Set of indices for effective samples
5:$\hat{A}_{\mathcal{E}}\leftarrow\{\}$triangleright Map: original index $i\in\mathcal{E}\to$ its absolute advantage $\hat{A}_{i}$
6:for $i\leftarrow 1$ to $N$ do
7:    $\hat{A}_{i,abs}\leftarrow|\tilde{A}_{i}|$triangleright Absolute advantage of current sequence $i$
@@
9:         $\mathcal{E}\leftarrow\mathcal{E}\cup\{i\}$triangleright Add index to effective set
10:         $\hat{A}_{\mathcal{E}}[i]\leftarrow\hat{A}_{i,abs}$triangleright Store absolute advantage for effective sample $i$
@@
15$\triangleright$ Step 3: Prioritized Sub-sampling from the Effective Set
16$\triangleright$ a. Calculate sampling probabilities $P_{j}$ for each $j\in\mathcal{E}$
17:    $Z\leftarrow\displaystyle\sum_{j\in\mathcal{E}}\exp(\hat{A}_{\mathcal{E}}[j]/\tau)$triangleright Normalization term (Softmax denominator over $\mathcal{E}$)
18:    $P_{\mathrm{dist}}\leftarrow\{\}$triangleright Map: original index $j\in\mathcal{E}\to$ its sampling probability $P_{j}$
@@
23:             $P_{\mathrm{dist}}[j]\leftarrow 1/|\mathcal{E}|$triangleright Uniform fallback if $Z=0$
@@
26$\triangleright$ b. Determine actual sub-sample size $k^{\prime}$
@@
28$\triangleright$ c. Sample $k^{\prime}$ indices from $\mathcal{E}$ according to probabilities $P_{\mathrm{dist}}$
29:    $S_{\mathrm{sampled\_indices}}\leftarrow\textsc{Sample}(\mathcal{E},\,P_{\mathrm{dist}},\,k^{\prime})$triangleright $S_{\mathrm{sampled\_indices}}$ is a list of $k^{\prime}$ indices from $\mathcal{E}$
30$\triangleright$ d. Form the distilled mini-batch
@@
A.3 Performance on Multimodal General Benchmarks
In this section, we provide an analysis in Table 6 to reveal that ReVisual-R1 demonstrates strong and competitive performance across these general MLLM benchmarks. Specifically, on the MMMU benchmark [84], Revisual-R1 secures the second-best score (50.55), closely following ThinkLite-7B-VL. It achieves leading performance on MM-Vet [85] with a top score of 49.81. In the MME-RealWorld benchmark [86], Revisual-R1 (62.68) delivers a solid performance, though it is surpassed by mmr1-mathv0 and ThinkLite-7B-VL. These results underscore Revisual-R1’s robust and well-rounded reasoning capabilities on these general multimodal tasks, particularly its notable strength on MM-Vet, within a competitive field where no single open-source 7B model consistently dominates all evaluated benchmarks.
Table 6: Performance of open-source multimodal reasoning 7B models on general MLLM benchmarks
Model MMMU MM-Vet MME-RealWorld
VL-Rethinker 42.22 47.59 61.43
OpenVLThinker 19.37 40.73 13.46
MMR1-Math-v0 47.33 42.75 69.12
R1-VL-7B 48.23 45.64 59.31
ThinkLite-7B-VL 51.22 37.44 68.49
VLAA-Thinker-7B 50.44 41.46 64.44
Revisual-R1 50.55 49.81 66.68
A.4 Reasoning Example

### v2

@@ 6.1 Multimodal Large Language Model
Multimodal Large Language Model (MLLM) is a key research area. While leading closed-source models (e.g., GPT-o3 (OpenAI, 2025), Kimi-VL (Kimi Team, 2025)) excel at long Chain-of-Thought (CoT) reasoning, open-source contributions have focused on CoT adaptations (Liu et al., 2023; Guo et al., 2024; Su et al., 2024) and Supervised Fine-Tuning (SFT) with reasoning traces (Mitra et al., 2024b; Zhang et al., 2024c) (Gao et al., 2024). DeepSeek-R1 (Guo et al., 2025a) has further spurred Reinforcement Learning (RL) applications for visual reasoning (Huang et al., 2025b) (Dong et al., 2024) and specialized domains like mathematical reasoning. Nevertheless, many MLLM reasoning models (Wang et al., 2025) (Deng et al., 2025) (Yang et al., 2025) (Chen et al., 2025b) (Huang et al., 2025a) are limited by generating relatively short responses, which often curtail genuine reflection, thorough visual exploration, and consequently, deep multimodal reasoning. Our work, in contrast, introduces a novel framework to enable MLLMs to generate significantly longer, reflective responses, thereby facilitating long CoT reasoning to unlock more comprehensive multimodal reasoning capabilities.
@@ 6.2 Reinforcement Learning in Reasoning
Current LLM research explores direct RL fine-tuning, specialized cold-start datasets for long-form reasoning, and advanced algorithms like Group Relative Policy Optimization (GRPO) (Shao et al., 2024) and its refinements (e.g., DAPO (Yu et al., 2025), DR.GRPO (Liu et al., 2025), GPG (Chu et al., 2025), AEPO (Chen et al., 2025c)) to elicit deeper reasoning. However, RL application to multimodal reasoning in MLLMs is nascent. Initial MLLMs efforts focus on subdomains like math reasoning (Meng et al., 2025b; Peng et al., 2025) or generative reward models (Gao et al., 2025), often utilizing data from commercial models. Nonetheless, successes such as DeepSeek-R1’s (Guo et al., 2025a) rule-based RL are spurring similar MLLMs investigations, indicating growing interest in RL for unlocking sophisticated multimodal reasoning.
6.3 Limitations and Future Work
Despite the strong empirical gains demonstrated by ReVisual-R1, our study has several limitations that point toward crucial directions for future work. While we justify the efficacy of the optimized cold start, staged multimodal and text-only RL, and PAD through substantial empirical gains and ablations, we currently lack a rigorous theoretical foundation. Specifically, a deeper theoretical account is needed to explain why these text-centric optimization strategies, such as the cold start, effectively enhance model reasoning capabilities, and to provide a more principled explanation for the staged RL methodology. Furthermore, our scalability evidence is limited to mid-sized models. We demonstrated that training on complex, high-difficulty textual reasoning data transfers benefits to both the textual and multimodal reasoning performance of 3B and 7B models. However, we have not yet validated these advantages across larger-scale architectures, such as MoE models, or generalized them to other foundational model families and larger size regimes. Finally, we will systematically investigate the interplay between data type and training stage. We will extensively explore the impact of the ratio and quality of multimodal versus textual data utilized during both the cold start and RL phases. The ultimate goal of this exploration is to determine an optimal curriculum recipe that jointly maximizes performance across diverse skill clusters, notably in STEM reasoning, perceptual grounding, and general domains.
@@ 7 Conclusion
This paper introduces ReVisual-R1, a 3B and 7B open-source MLLM designed to address prevalent challenges in cultivating sophisticated multimodal reasoning. By systematically integrating a strategic, high-difficulty text-only cold-start phase for foundational reasoning, a Multimodal RL stage employing GRPO stabilized by our novel Prioritized Advantage Distillation (PAD) mechanism, and a final TextRL refinement phase, our structured three-stage curriculum demonstrates that thoughtful data strategy and targeted algorithmic optimizations are pivotal. ReVisual-R1 achieves superior performance among open-source 7B models on a suite of challenging multimodal, textual reasoning and generalization benchmarks. This work underscores that careful curriculum design and algorithmic enhancements, rather than sheer model scale, can unlock robust, self-reflective multimodal reasoning.
Reproducibility Statement
To ensure the reproducibility of our work, we provide an anonymous code repository in the supplementary materials containing the full implementation of our Revisual-R1 algorithm. All datasets used for training and evaluation are publicly available, and a comprehensive breakdown of our experimental setup, including all key hyperparameters, is detailed in Appendix.
Ethics Statement
The development of our model, ReVisual-R1, and the curation of the GRAMMAR dataset were conducted with close attention to ethical guidelines. Our research is grounded in the use of publicly available and open-source academic datasets, which were rigorously filtered during the creation of GRAMMAR to ensure no personal, private, or sensitive information was included. The work does not involve human subjects, and its objective is to advance research in multimodal reasoning for the open-source community. We foresee no direct societal risks or harmful applications.
Appendix
Appendix Contents
Appendix A LLM Usage Statement
In this paper, we employed a Large Language Model (LLM) in an assistive capacity for this manuscript. Its use was strictly limited for grammatical accuracy and refining sentence structure to improve clarity and readability. All intellectual contributions, including the formulation of ideas, data analysis, and the composition of the manuscript, are entirely the work of the authors.
Appendix B Training settings
@@
Table 6: Performance comparison on general textual and multimodal benchmarks. Best results per group (3B or 7B) are bold, and second-best per group (3B or 7B) are underlined.
Model MMMU MMMU-PRO CMMMU MMStar MMLU-Pro Avg. 3B Models Qwen2.5-VL-3B-Instruct $49.78$ $32.64$ 45.93 $51.61$ $41.24$ $44.64$ FAST-3B 52.00 34.61 $44.58$ 55.15 44.06 46.48 VLAA-Thinker-3B $49.67$ $33.34$ $44.92$ $53.35$ $42.35$ $44.33$ Revisual-R1-3B 64.78 44.47 51.19 53.76 51.82 53.60 7B Models Qwen2.5-VL-7B-Instruct $52.00$ $42.52$ $51.69$ $61.71$ $51.80$ $51.94$ ThinkLite-7B-VL $53.00$ $41.01$ $50.51$ 64.73 $52.74$ $52.34$ VL-Rethinker-7B 59.78 45.59 52.54 $63.92$ $53.31$ 55.03 MM-Eureka-Qwen-7B $56.22$ $43.50$ $52.20$ 64.16 54.03 $54.17$ Revisual-R1-7B 67.00 54.35 56.95 $63.86$ 62.38 60.91
Appendix C Algorithm in Prioritized Advantage Distillation (PAD)
@@
Initially, PAD filters the original batch $\mathcal{B}$ to create an “effective set” $\mathcal{E}$ of sample indices and a corresponding map $\hat{A}_{\mathcal{E}}$ for their advantages (Lines 2-10 in Algorithm 1). For each sequence $i$ in $\mathcal{B}$, its absolute advantage $|\tilde{A}_{i}|$ is computed. If this value falls within a specified informative range $T_{low},T_{high}]$, where $T_{low}>0$ is crucial for excluding stagnant (near-zero advantage) samples, the index $i$ is added to $\mathcal{E}$, and its absolute advantage $\hat{A}_{i,abs}$ is stored in $\hat{A}_{\mathcal{E}}$.
@@
1:Original batch $\mathcal{B}=\{seq_{1},\dots,seq_{N}\}$ with $N=|\mathcal{B}|$; advantage estimates $A_{\mathrm{est}}=\{\tilde{A}_{1},\dots,\tilde{A}_{N}\}$; thresholds $T_{\mathrm{low}},\,T_{\mathrm{high}}$; temperature $\tau$; sub–sampling ratio $\rho$
2:Distilled mini‑batch $\mathcal{B}_{\mathrm{distilled}}$
[…]
6:for $i\leftarrow 1$ to NN do
7:  $\hat{A}_{i,abs}\leftarrow|\tilde{A}_{i}|$ $\triangleright$ Absolute advantage of current sequence $i$
@@
9:   $\mathcal{E}\leftarrow\mathcal{E}\cup\{i\}$ $\triangleright$ Add index to effective set
10:   $\hat{A}_{\mathcal{E}}[i]\leftarrow\hat{A}_{i,abs}$ $\triangleright$ Store absolute advantage for effective sample $i$
@@
[…]
17:  $Z\leftarrow\displaystyle\sum_{j\in\mathcal{E}}\exp(\hat{A}_{\mathcal{E}}[j]/\tau)$ $\triangleright$ Normalization term (Softmax denominator over $\mathcal{E}$)
18:  $P_{\mathrm{dist}}\leftarrow\{\}$ $\triangleright$ Map: original index $j\in\mathcal{E}\to$ its sampling probability $P_{j}$
@@
23:     $P_{\mathrm{dist}}[j]\leftarrow 1/|\mathcal{E}|$ $\triangleright$ Uniform fallback if $Z=0$
@@
[…]
@@
[…]
29:  $S_{\mathrm{sampled\_indices}}\leftarrow\textsc{Sample}(\mathcal{E},\,P_{\mathrm{dist}},\,k^{\prime})$ $\triangleright$ $S_{\mathrm{sampled\_indices}}$ is a list of $k^{\prime}$ indices from $\mathcal{E}$
[…]
@@
Appendix D Reasoning Example